OpenStudy (kingpop):

if I have a 2x + 6=5x + 8 what do I do

OpenStudy (otfgloboy):

add ur 2x to ur 5x

OpenStudy (anonymous):

no that would be wrong

OpenStudy (otfgloboy):

or u could subtract it to

alones (alones):

Hey there Welcome to Openstudy =] 2x+6−5x=5x+8−5x −3x+6=8 Don't forget to ubtract 6 from both sides −3x+6−6=8−6 −3x=2 Step 3: Divide both sides by -3. −3x/−3=2/−3 and you should get x=−2/3

myininaya (myininaya):

Let's try to get the asker involved just a little bit more next time @AloneS . Thanks.

OpenStudy (anonymous):

2x+6=5x+8 -5 -5 ---------- -3x+6=8 -6 -6 ---------- -6x=8 divide by -6 on both sides ---------------------- x=-1.333333333333333333333

alones (alones):

Oh my bad xP thanks @myininaya I'll try =]

OpenStudy (kingpop):

thanks everyone

OpenStudy (otfgloboy):

is that bftd 3 by chief keef bro